100ml of a liquid contained in an insulated container at a pressure of 1 bar. The pressure is steeply increasedto 100 bar. The volume of the liquid is decreased by 1ml at this constant pressure. Find the ΔH & ΔU....

Grade:12

2 Answers

vijesh jayan
13 years ago

V1=100ml & V2=?,

P1=1bar & P2=100bar

ΔH=ΔU+nRT

BY CHARLES LAW

P1V1=P2V2

1*100=100*V2

100/100=V2

V2=1ml

IT IS AN INSULATED CONTAINER SO NO TRANSFER OF HEAT TAKES PLACE

THEREFORE ΔU=0

ΔH=0+nRT  =>  ΔH=nRT

PV=nRT [IDEAL GAS EQUATION]

(P1-P2)*(V1-V2)=(1-100)*(100-1)=-99*99=-9801

IS WHAT I THINK SO

I AM NOT SURE

ENQUIRE OR ASK SOME ONE ELSE

Aiswarya Ram Gupta
13 years ago

thnx...............